TWO DIFFERENT SOLUTIONS OF ALCOHOL WITH RESPECTIVE PROPORTIONS OF WATER TO ALCOHOL OF 3:1 AND 3:2 WERE COMBINED .WHAT IS THE CONCENTRATION OF ALCOHOLIN THE NEW SOLUTION SOLUTION IF THE ORIGINAL SOLUTIONS WERE MIXED IN EQUAL AMOUNT
PLZ Help me with this one
best regards
mandy

Answer by checkley71(8403)

1/3+2/3=X OR 3/3=2X 1=2X OR X=1/2
OR 1.2*1/3+1/2*2/3=X OR 1/6+1/3=X OR 1/6+2/6=X OR 3/6=X OR X=1/2